# FedDRL: A Trustworthy Federated Learning Model Fusion Method Based on Staged Reinforcement Learning

### 21 Jan 2024 v3

@@ Abstract
Traditional federated learning uses the number of samples to calculate the weights of each client model and uses this fixed weight value to fusion the global model. However, in practical scenarios, each client’s device and data heterogeneity leads to differences in the quality of each client’s model. Thus the contribution to the global model is not wholly determined by the sample size. In addition, if clients intentionally upload low-quality or malicious models, using these models for aggregation will lead to a severe decrease in global model accuracy. Traditional federated learning algorithms do not address these issues. To solve this probelm, we propose FedDRL, a model fusion approach using reinforcement learning based on a two staged approach. In the first stage, Our method could filter out malicious models and selects trusted client models to participate in the model fusion. In the second stage, the FedDRL algorithm adaptively adjusts the weights of the trusted client models and aggregates the optimal global model. We also define five model fusion scenarios and compare our method with two baseline algorithms in those scenarios. The experimental results show that our algorithm has higher reliability than other algorithms while maintaining accuracy.
@@ 1 Introduction
With the development of deep learning technologies, numerous industries are applying them to their respective fields, resulting in the emergence of intelligent transportation, smart logistics, and smart healthcare models. These technologies reduce the industry’s production and management costs, increase operational efficiency, and accelerate the industry’s digitization process. However, most current deep learning-based model training uses a supervised learning method. The number and diversity of samples is the key to training a high-quality model. Therefore, how to obtain massive and diverse data samples has become the first step to training deep learning models. This method leads to the source of these samples being no longer limited to a single industry but to multiple industries collaborating to construct large-scale datasets. Moreover, with the emergence of cross-industry collaborative data analysis models, data analysis sources are no longer limited to a single organization but involve multiple parties. Although the participation of multiple parties in the data analysis process can solve the problem of cross-domain data analysis, this strategy also faces a new challenge. The traditional data analysis method requires uploading each participant’s raw data to a central node and training models on the central node. However, this workflow involving raw data transfer poses security and privacy risks. Meanwhile, raw data transmission is not permitted for data-sensitive industries, such as hospital patient and financial industry data. How to achieve federated data analysis of multiple parties while protecting data security is becoming hot research.
Federated learning is an emerging paradigm for distributed computing. The structure of federated learning has a central node and multiple participants, where multiple participants train models with their private data, and the central node achieves knowledge extraction and information fusion of each participant by fusing the models uploaded by each client. Since the process does not directly use the participants’ data, federated learning can achieve joint data analysis of multiple participants while protecting user privacy in advance. To solve mobile device word input privacy protection issues, Google proposed federated learning and the FedAvg method [1], which ensure user privacy by transmitting models instead of raw data. Increasingly, companies are applying federated learning in various domains to tackle cross-industry data analysis challenges. However, in practical federated learning environments, the heterogeneity of data among participants gives rise to the non-iid problem, which makes model convergence challenging during the fusion stage of the FedAvg method. Additionally, variations in computational resources among participants lead to slow convergence of the global model. To address these challenges, some research has optimized the global model convergence problem by adding control parameters in the client-side model fusion process, such as FedProx [2], Scaffold [3], and FedNova [4]. The Moon algorithm [5] optimizes the convergence of the global model by introducing a comparative loss between the client-side model and the global model. There is also some research [6] [7] to improve the accuracy of the global model by selecting the clients involved in the global model. These algorithms somewhat solve the global model’s convergence problem in the non-iid data among participants.
Current traditional federal learning methods, such as FedAvg and FedProx, use a ratio based on the number of samples of each participant to the total number of samples to calculate the model weights of each participant. However, in a practical scenario, the method has two problems. On the one hand: the method does not consider the different participants’ model quality, which gives higher weights to the client with a larger number of samples. However, due to the variability of computational resources and sample quality across clients, it does not mean that a client can train a higher-quality model when a client has more samples. Therefore, using these algorithms assumes that the sample quality and computational resources are consistent across clients. On the other hand, if the participants have malicious attack behavior and intentionally upload malicious models to participate in fusion, which will seriously influence the global model convergence process. Therefore it becomes increasingly important to design adaptive defenses for federated learning systems [8]. How to eliminate malicious models in federated learning model fusion has become a vital issue. At the same time, when the client uploads low-quality models to participate in fusion, how to adaptively adjust each model’s fusion weights based on the models’ quality is also an urgent problem to be solved. Some research has used reinforcement learning methods to address these issues of weighting. Such as, Favor [7] method uses the DDPG to calculate the weights of the participant models. Some research has also used reinforcement learning to solve the problems of device selection [9] [10], resource optimization [11] [12], and communication optimization in IoT federated learning scenarios.
Reinforcement learning (RL) methods use a try-and-learn strategy. The core of the reinforcement learning approach is to train an intelligent agent. The agent uses different actions to interact with the external environment. The environment provides reward and punishment values based on the selected actions of the intelligence. The intelligent body achieves optimal action selection by maximizing the reward value. However, when we use reinforcement learning, we must address some issues. On the one hand, we must collect training samples by continuously interacting with the environment. When the cost of interacting with the external environment is too high, or the interaction cost is unacceptable (e.g., driverless scenario), how to collect samples quickly will directly affect the training time of reinforcement learning. On the other hand, when the action space of the agent is high-latitude and continuous, this will lead to an increasing sampling time. Therefore, due to these two problems, traditional single-agent reinforcement learning training methods can take much time. When we apply reinforcement learning to federated learning, we must solve those problems. As the number of participants grows, the time cost of training agents will increase. How to optimize the action space for reinforcement learning and speed up the training process of the agent is an urgent problem to be solved.
To address these issues, we propose a staged reinforcement learning algorithm (FedDRL) to solve the problem of model weight calculation in federated learning. The algorithm consists of two phases: the selection of trusted clients and the adaptive assignment of weights. We use a staged approach to accelerate the agent’s training. We use a distributed reinforcement learning algorithm to speed up the collection of training samples. The contributions of this paper is as follows.
  • We propose a federated learning model fusion method based on reinforcement learning, which solves the problem of malicious clients’ participation in model fusion and clients’ model weight assignment.
  • We propose a trusted client selection strategy that eliminates malicious models from participating in global model fusion while enabling the selection of trusted clients.
@@
  • We propose a adaptive weight assignment method that adaptively adjusts the weights corresponding to the clients according to the quality of their uploaded models and therefore computes the optimal global model.
@@ 2.1 Federated learning
Research of federated learning focuses on solving three problems: how to improve generalization of the global model on the server side, how to personalize the model on the client side.Therefore, Federated learning can be divided into server-side and client-side optimization algorithms. Google first proposed the FedAvg method [1] to solve the server-side global model fusion problem . However, the client in the actual scenario has device and sample heterogeneity problems, which can lead to global model convergence problems in the FedAvg method. To optimize the global model convergence problem, Karimireddy et al. proposed the Scaffold algorithm [3], which accelerates the convergence of the global model by introducing a control variable to correct the client-side drift problem. Li et al. proposed the FedProx algorithm [2], which implements the correction of deviated clients by adding the regularization function to the client model. Wang et al. proposed the FedNova algorithm [4], which solves the global model convergence problem by normalizing the parameters on both the client and server sides. Li et al. also proposed the MOON [5] method based on model comparison learning, which solves the global model convergence problem by calculating the similarity between the global model and the client model and continuously decreasing the distance representation of both to achieve the training of the client-side model.
2.2 The challenges of Non-IID Data Distribution
The Non-IID problem is the key that leads to the convergence of the federated learning model. Zhao et al. study the performance of different federated learning methods on the Non-IID dataset and experimentally demonstrate that the Non-IID problem poses a significant challenge to the accuracy of the federated learning algorithm [13]. Therefore some research focuses on solving the Non-IID problem of federated learning. For example, Zhang et al. proposed the FedPD method [14], which offers strategies that can handle non-convex objective functions to optimize models and communication. In addition, Gong et al. proposed the AutoCFL method [15], which adaptively adjusts the number of training rounds for the client while introducing weighted voting of the client clustering strategy to eliminate the effects of Non-IID and imbalanced data. Huang et al. proposed the FedAMP method [16], which solves the client-side model personalization problem caused by Non-IID data by implementing a personalized update of the client-side model. Li et al. proposed the Fedbn method [17], which adds a batch normalization layer to the local model to solve the problem of feature shift in data heterogeneity. Briggs et al. proposed a hierarchical clustering approach (FL+HC)[18], which optimizes the model’s performance on Non-IID datasets by training clients in groups and then having each group trained independently on a dedicated model. Furthermore, Gao et al. proposed the Feddc method [19], which addresses the data heterogeneity problem by bridging the gap between client and global model parameters through a control variable. Finally, mu et al. proposed the Fedproc method [20], which guides the client model training by introducing a loss of comparison between the client model and the global model, thus enabling the global model to perform well on Non-IID data to achieve good performance.
@@ 2.3 Federated reinforcement learning
Since reinforcement learning has the advantage of adaptive learning, some research applies reinforcement learning to federated learning. Some research focuses on improving the performance of global models based on reinforcement learning methods. Such as Wang et al. proposed the Favor method [7], which adaptively selects clients participating in the model fusion. Sun et al. also proposed the PG-FFL framework [6], which can solve the problem of calculating client weights during model fusion. Some work uses reinforcement learning to solve device optimization problems in a federated IoT environment. such as Zhang et al. uses the DDPG method [9] to achieve the selection of optimal devices. Zhang proposed the FedMarl algorithm [21], which uses Multi-agent reinforcement learning to achieve node selection in federated learning. Yang et al. also proposed a digital twin architecture (DTEI) [10], which uses reinforcement learning methods to solve the device selection problem. Some research focus on solving the resource optimization and scheduling problem in IoT environment. Zhang et al. proposed the RoF approach[11], which achieves optimal resource scheduling based on multi-intelligent reinforcement learning. Rjoub et al. designed a trusted device selection method [22] based on a reinforcement learning approach to achieve resource scheduling. Rjoub et al. proposed the DDQN-Trust method [12], which uses the Q-learning approach to evaluate the credit scores of devices to achieve optimal scheduling. To optimize the communication problem of federated learning, Yang et al. proposed a reinforcement learning-based model evaluation method [23] to participate in model training and model fusion by selecting the optimal devices. Most current research uses reinforcement learning to solve federated learning problems in IoT environments. These include optimal device selection, resource optimization, and communication optimization. Some work uses reinforcement learning methods to solve the client-side weight assignment problem in federated learning. However, most of these methods are based on single-agent training methods and need to address the training time overhead of reinforcement learning.
@@ 3.1 Problem Definition
In this section, we first analyze the problems of the current federated learning approach and then propose a solution to the problem. In the federation learning process, we aim to get a higher-accuracy global model by aggregating all clients’ local models through the server side.
In the federated learning environment, we define n clients are involved in model fusion, and the client is denoted as $C_{i}$ where $C_{i}\in\left\{C_{1},C_{2},C_{3}\ldots C_{n}\right\}$. Each client has a network model $M_{i}$, where $M_{i}\in\left\{M_{1},M_{2},M_{3}\ldots M_{n}\right\}$. Each client has its private data $D_{i}$, where $D_{i}\in\left\{D_{1},D_{2},D_{3}\ldots D_{n}\right\}$. The number of samples in each dataset is $S_{i}$, where $S_{i}\in\left\{S_{1},S_{2},S_{3}\ldots S_{n}\right\}$. The total number of samples is $\sum_{i=1}^{N}S_{i}$. We define the $\theta_{i}$ as a model parameter of $M_{i}$. where $\theta_{i}\in\left\{\theta_{1},\theta_{2},\theta_{3}\ldots\theta_{n}\right\}$.
Also, we define the server-side model aggregation process for each round as equation 1.
@@
Where $w_{i}$ is the weight of each client model parameter. The traditional method of calculating this weight in federal learning is determined by the ratio of the number of samples of the client to the total number of samples. This means that the larger the number of samples corresponding to the client, the larger the weight assigned. However, in the actual environment, due to the heterogeneity of the client data and the computing resources, it is impossible to determine the weights by the ratio of the number of samples for each client. Therefore, we need to solve two problems.
Problem 1: The sample size of a client accounts for 20% of the total samples, but the accuracy of this client model is only 53%, which will reduce the accuracy of the global model if the traditional federated model fusion algorithm is used.
Problem 2: If the client involved in the model fusion has malicious attacks and intentionally modifies the model output to make its model accuracy 10%, the global model accuracy will be seriously affected if the malicious models are selected for fusion using the traditional federated model fusion algorithm.
To solve these two problems, we need an adaptive weight calculation method that can set the weight of the malicious model to 0, which means that malicious models are prohibited from participating in the fusion process. At the same time, we realize the adaptive assignment of weights to each client and try to increase the weights of high-quality client models so that the global model gets the optimal accuracy rate.
If we use the single-agent reinforcement learning approach to solve the above problems, we face another issue again. On the one hand, the action space of the agent can be huge as the number of fused clients increases, which also leads to a long time required to train the agent model. On the other hand, a single agent method can only interact with a single environment, and that sampling process takes a long time. We solve these two problems in two steps by borrowing the idea of hierarchical reinforcement learning, which can solve the problem of the long training time of reinforcement learning. These two stages are selecting trusted clients and the optimal weight assignment process.
Stage 1: In this phase, We need select K-trusted models from N models to participate in the global model fusion. However, we do not know which client uploaded a malicious model. To solve the problem, We use reinforcement learning to achieve dynamic selection and autonomous screening of client models as equation 2.
$\left\{M_{a},M_{\mathrm{b}},\ldots M_{\mathrm{k}}\right\}\leftarrow\text{ SelectTrustedModel }\left(\left\{M_{1},M_{2},\ldots M_{\mathrm{n}}\right\}\right)$ (2)
@@
Stage 2: Based on step 1, we assign optimal weights to the trusted models to improve the accuracy of the global model and represent the process as equation 3.
$\left\{W_{1},W_{2},\ldots W_{\mathrm{n}}\right\}\leftarrow\operatorname{AdaptCalculateWeight}\left(\left\{M_{a},M_{b},\ldots M_{\mathrm{k}}\right\}\right)$ (3)
@@
$AdaptCalculateWeight(.)$ is a weight-adaptive computation method. $W_{i}$ is the optimal computational weight for each client of the output.
Why do we use a phased approach to reinforce learning? Let’s use an example to illustrate the reason for using it. We need a robot to learn the cooking process using a reinforcement learning approach. The cooking steps are divided into three stages: washing, chopping, and cooking. The robot needs to learn the best actions to cook a good dish. Traditional reinforcement learning is to find the best action that satisfies all three stages at the same time. However, if the robot does not wash the vegetables well, it cannot cook a good dish in the subsequent stages either. Therefore, it is a complete waste of time to try the next step of action if the optimal action of the previous stage is not learned. If the robot learns in multiple stages, it first learns the optimal action in the cleaning process and then proceeds to the next step of learning the optimal action. This will save time and make it easier to get the best solution.
Therefore, if the malicious models are not screened and the weight assignment method is used for malicious models, it will increase the cost of trial and error time in the reinforcement learning process. Therefore, we screen the trusted clients for weight assignment by a staged approach. This can exclude the trial and error cost of malicious models involved in weight assignment, thus optimizing the action space. Also, we use distributed reinforcement learning methods to increase the sampling speed and thus reduce the time overhead of training reinforcement learning. We will describe the implementation of each step in detail in the next section.
3.2 Trusted Client Selection Method
We propose the FedDRL approach to solve the two problems in the previous section. In the first stage, we select the trusted clients to achieve the fusion of the models by excluding the malicious clients. In the second stage, we compute the optimal model weights to improve the accuracy of the global model. The flow of the FedDRL method is shown in Figure 1.
We select trusted clients using the A2C algorithm. A2C is a synchronization algorithm achieved by improving the A3C algorithm [24]. The selection process of a trusted client for each communication round of federated learning can be treated as a Markov decision process. The reinforcement learning modeling in this phase is described as follows.
State: For the m-th communication round, we collect n models uploaded by each client and select k clients to participate in the fusion of global models. We define the $S^{t}$ as a select k client to fusion model in state t, which is shown as Equation 4, and the state includes the selected client models, the accuracy of each model, and the accuracy of the global model aggregated out by the selected client.
$S^{t}=\left\{\theta_{1}^{t},\theta_{2}^{t},\theta_{3}^{t}\ldots\theta_{k}^{t},acc_{1}^{t},acc_{2}^{t},acc_{3}^{t}\ldots acc_{k}^{t},acc_{global}^{t}\right\}$ (4)
Action: Our main objective is to eliminate the malicious models and select the trusted models to participate in the fusion. To achieve this purpose, we randomly choose several models to fusion the global model and evaluate the accuracy of global models. If a malicious model is selected to participate in the fusion, the accuracy of the global model will be very low. If only trusted clients are selected to participate in the fusion, the accuracy of the global model will also increase. We define the action of selecting a model as $a^{t}\in\left\{0,1\right\}$, where one means that the model is selected and 0 means that the model is not selected. Thus, n models correspond to the action space that can be expressed as equation 5.
$A^{t}=\left\{a_{1}^{t},a_{2}^{t},a_{3}^{t}\ldots a_{n}^{t}\right\},\quad a^{t}\in\left\{0,1\right\}$ (5)
Reward: We define the accuracy of selecting all models aggregated as $Acc_{all}$; we take this accuracy as the baseline and also represent the global model accuracy $Acc_{m}$ obtained by the m-th randomly selected client models, and calculate the reward value $Reward_{1}$ by the difference of $Acc_{m}-Acc_{all}$. Meanwhile, to avoid the intelligent body only, It can be realized that the agent selects all trusted clients as much as possible. The higher the number of models the agent sets, the higher the value of this reward. Therefore, the total reward is defined as equation 6.
$Reward=\begin{cases}\alpha*(Acc_{m}-Acc_{all})+\beta*(Number),Acc_{m}>Acc_{all}\\ \alpha*(Acc_{m}-Acc_{all}),Acc_{m}<Acc_{all}\end{cases}$ (6)
Where $\alpha$ and $\beta$ are fixed values that can be custom adjusted to balance the importance of Reward1 and Reward2, preventing one of the rewards from being too large and affecting agent training.
The agent’s goal is to obtain the maximum long-term reward value based on the discount factor $\gamma$, and the process is expressed as equation 7.
$R(\tau)=\sum_{\mathrm{t}=1}^{\mathrm{T}}\gamma^{t}\mathrm{r}_{\mathrm{t}}$ (7)

### 19 Mar 2024 v4

@@ Abstract
Federated learning facilitates collaborative data analysis among multiple participants while preserving user privacy. However, conventional federated learning approaches, typically employing weighted average techniques for model fusion, confront two significant challenges: (1) The inclusion of malicious models in the fusion process can drastically undermine the accuracy of the aggregated global model. (2) Due to the heterogeneity problem of devices and data, the number of client samples does not determine the weight value of the model. To solve those challenge, we propose a trustworthy model fusion method based on reinforcement learning (FedDRL), which includes two stages. In the first stage, we propose a reliable client selection mechanism to exclude malicious models from the fusion process. In the second stage, we propose an adaptive model fusion method that dynamically assigns weights based on model quality to aggregate the best global models. Finally, We validate our approach against five distinct model fusion scenarios, demonstrating that our algorithm significantly enhances reliability without compromising accuracy.
@@ 1 Introduction
With the advent of deep learning technologies, various industries have been integrating these technologies into their sectors, promoting the development of intelligent transportation, smart logistics, and healthcare systems. These technologies are crucial in reducing production and management costs, enhancing operational efficiency, and accelerating industry digitization. However, supervised learning remains the primary method for training deep learning models, where the volume and diversity of samples are essential for creating high-quality models. Consequently, acquiring extensive and varied data samples has emerged as the initial step in training deep learning models. This approach has led to sample sources expanding from single industries to collaborations across multiple sectors to develop large-scale datasets. To achieve multi-party joint data analysis under the condition of protecting data security and privacy, Google has proposed federated learning technology for the first time. Although federated learning solves the problem of user privacy protection, the traditional federated learning algorithm assumes that all participants are trustworthy. On the contrary, in the actual scenario, if participants exhibit malicious behavior and intentionally contribute harmful models to the fusion process, it can significantly disrupt the global model’s convergence. Thus, creating adaptive defenses for federated learning systems becomes increasingly crucial [9]. Identifying methods to remove malicious models in federated learning model fusion has become a critical issue. Simultaneously, when a client submits low-quality models for fusion, determining how to adaptively adjust each model’s fusion weights based on their quality is also an urgent problem needing resolution. Some studies have applied reinforcement learning techniques to address these weighting issues. For instance, the Favor [10] method uses the DDPG to assign weights to participant models. Additional research has applied reinforcement learning to address device selection [11] [12], resource optimization [13] [14], and communication optimization in IoT federated learning contexts.
Reinforcement learning (RL) employs a trial-and-error strategy. The essence of this approach is training an intelligent agent that interacts with the external environment through varied actions. The environment then provides feedback in the form of rewards and penalties based on the agent’s actions, guiding the agent toward optimal action selection by maximizing reward value. However, employing reinforcement learning presents certain challenges. Firstly, continuous training is required for sample collection through environmental interaction. When the cost of such interactions is prohibitive or unacceptable (for example, in our scenario, where the server must frequently calculate the global model’s parameters), the efficiency of sample collection significantly impacts the reinforcement learning training duration. Secondly, when the agent’s action space is vast and continuous, it leads to prolonged sampling periods. These issues mean traditional single-agent reinforcement learning training approaches can be exceedingly time-consuming. Applying reinforcement learning in federated learning requires addressing these problems, as increasing participant numbers escalates agent training time. Therefore, optimizing the action space for reinforcement learning to expedite the agent training process is an essential challenge to address.
Why opt for phased reinforcement learning? We take an example to explain this problem. Consider a robot learning to cook through reinforcement learning, with the process divided into washing, chopping, and cooking stages. The robot must master each stage to prepare a successful dish. Traditional reinforcement learning aims to identify the optimal action across all stages simultaneously; however, mastering the initial stage is essential before progressing. By adopting a phased learning approach, the robot sequentially masters each stage, streamlining the learning process and leading to more effective outcomes. Similarly, if malicious models are not initially filtered out, the agent’s trial-and-error costs in weight assignment for these models will increase. To resolve these issues, we propose a staged reinforcement learning algorithm (FedDRL). The contributions of the paper are as follows.
  • We design a federated learning framework that employs reinforcement learning for model fusion, designed to select trustworthy clients and optimally assign model weights.
  • We propose an adaptive client selection strategy based on the A2C algorithm, dynamically identifying and selecting trustworthy clients while excluding malicious ones from the model fusion process based on situational analysis.
@@
  • We propose an adaptive weight assignment method that adaptively adjusts the weights according to the quality of their uploaded models.
  • We propose an adaptive weight assignment method that adaptively adjusts the model fusion weights according to the quality of their uploaded models.
@@ 2.1 Federated Learning
Research in federated learning primarily aims to address two challenges: enhancing the generalization of the global model on the server side and personalizing the model on the client side. Consequently, federated learning algorithms are bifurcated into server-side and client-side optimization strategies. Google initially introduced the FedAvg algorithm [2] to address the problem of server-side global model fusion. To improve global model convergence, Karimireddy et al. developed the Scaffold method [3], which mitigates client-side drift by integrating a control variable. Similarly, Li et al. introduced FedProx [4], applying a regularization function to client models to correct deviations. Additionally, Wang et al. unveiled FedNova [5], addressing global model convergence issues by normalizing parameters on both client and server ends. Furthermore, Li et al. have introduced the MOON [6] technique, leveraging model comparison learning to enhance global model convergence. Chen et al. [29]. also proposed a client identification method based on model parameter features to achieve trustworthy federated learning.
While those approaches enhance the global model’s convergence speed, practical federated learning situations reveal variances in the quality of models trained by individual participants. These discrepancies stem from the diversity in computational resources and the calibre of data samples available to each participant. Additionally, variations arise due to the quantity and type of samples possessed by each participant, a phenomenon known as Non-IID (Non-Independent and Identically Distributed). Consequently, these factors complicate the attainment of optimal global model aggregation in the Non-IID environments.
2.2 Challenges of Non-IID Data Distribution
The Non-IID data issue significantly impacts federated learning models’ convergence. Zhao et al. explored various federated learning methods’ performance on non-IID datasets, demonstrating significant accuracy challenges [15]. Accordingly, several studies have addressed the non-IID dilemma in federated learning. For instance, Zhang et al. proposed the FedPD approach [16], optimizing models and communication for non-convex objective functions. Moreover, Gong et al. introduced AutoCFL [17], utilizing a weighted voting client clustering strategy to mitigate non-IID and imbalanced data effects. Huang et al. developed FedAMP [18], which addresses Non-IID data-induced client-side model personalization issues through personalized model updates. Li et al. devised Fedbn [19], incorporating a batch normalization layer into local models to address feature shift challenges due to data heterogeneity. Briggs et al. suggested a hierarchical clustering method (FL+HC) [1], improving Non-IID dataset model performance by grouping clients for independent model training. Additionally, Gao et al. offered the Feddc approach [20], bridging client and global model parameter disparities through a control variable. Lastly, Mu et al. introduced Fedproc [21], directing client model training by integrating a comparative loss between client and global models. Chen [7] et al. proposed a federated learning method based on adaptive knowledge distillation to improve the accuracy of heterogeneous model scenarios.
Although these methodologies advance Non-IID issue mitigation in federated learning, they typically assign uniform fusion weights to all clients, failing to exclude malicious or low-quality model contributions. Consequently, dynamically selecting clients for fusion and adaptively calculating each model’s weight remains critical for successful global model integration.
@@ 2.3 Federated Reinforcement Learning
Given the adaptive learning potential of reinforcement learning, its application within federated learning contexts has garnered interest. Some research has concentrated on leveraging reinforcement learning to boost global model performance. For instance, Wang et al. introduced the Favor method [10], which adaptively selects clients for model fusion. Sun et al. developed the PG-FFL framework [22], addressing the challenge of client weight computation during model fusion. Additional studies have applied reinforcement learning for device optimization within federated IoT frameworks. For example, Zhang et al. utilized the DDPG algorithm [11] for optimal device selection. Zhang also formulated the FedMarl strategy [23], employing multi-agent reinforcement learning for node selection. Similarly, Yang et al. proposed a digital twin architecture (DTEI) [12], applying reinforcement learning for device selection issues. Other investigations have addressed resource optimization and scheduling challenges within IoT contexts, such as Zhang et al.’s RoF methodology [13], which leverages multi-intelligent reinforcement learning for optimal resource scheduling. Additionally, Rjoub et al. have developed trusted device selection techniques [24] and the DDQN-Trust method [14], utilizing Q-learning to assess devices’ credit scores for optimal scheduling. To ameliorate federated learning communication issues, Yang et al. introduced a reinforcement learning-based model evaluation method [25], selecting optimal devices for training and fusion. Nevertheless, while these efforts predominantly focus on IoT environment applications—such as device selection, resource optimization, and communication enhancement—they seldom address federated learning’s model weight calculation challenges. Therefore, Zhang et al. proposed the $R^{2}$ed framework [27], employing the DDPG reinforcement learning method for adaptive client weight calculation. Chen et al. [28] designed a task platform for implementing trustworthy federation learning.
Although current research addresses the issue of weight allocation in federated learning, it often neglects the training efficiency of the agents. Therefore, optimizing the training efficiency of agents is a significant challenge that needs attention.
@@ 3.1 Problem Definition
In this section, we scrutinize the prevailing challenges of the current federated learning approach and subsequently propose a solution. In federated learning, the objective is to get the global model by amalgamating local models from all clients through server-side aggregation. We define n clients as involved in model fusion, and the client is denoted as $C_{i}$ where $C_{i}\in\left\{C_{1},C_{2},C_{3}\ldots C_{n}\right\}$. Each client has a network model $M_{i}$, where $M_{i}\in\left\{M_{1},M_{2},M_{3}\ldots M_{n}\right\}$. Each client has its private data $D_{i}$, where $D_{i}\in\left\{D_{1},D_{2},D_{3}\ldots D_{n}\right\}$. The number of samples in each dataset is $S_{i}$, where $S_{i}\in\left\{S_{1},S_{2},S_{3}\ldots S_{n}\right\}$. The total number of samples is $\sum_{i=1}^{N}S_{i}$. We define the $\theta_{i}$ as a model parameter of $M_{i}$. where $\theta_{i}\in\left\{\theta_{1},\theta_{2},\theta_{3}\ldots\theta_{n}\right\}$.
Additionally, the server-side model aggregation process per round is defined as shown in equation 1:
@@
The $w_{i}$ is the fusion weight of each model parameter.
Traditional Federated Learning typically employs a weighted average approach for computing model fusion weights, with each model’s weight determined by its corresponding client’s data sample size relative to the total. Thus, clients contributing more data exert a greater influence on the aggregated model. However, this method fails to consider the quality of each client’s model and the potential inclusion of malicious models in real-world scenarios. We illustrate the deficiencies of the traditional federated fusion algorithm through two scenarios:
Scenario 1: A client’s data represents 20% of the total, yet its model’s accuracy is merely 53%. Employing the conventional federated fusion algorithm in this case would detrimentally impact the global model’s accuracy.
Scenario 2: A client engaged in model fusion launches malicious attacks, intentionally skewing its model’s output to reflect a mere 10% accuracy. If such malicious models are incorporated through the standard fusion process, the accuracy of the global model would be severely compromised.
Addressing these challenges necessitates an adaptive weight calculation strategy capable of nullifying malicious models by assigning them a weight of zero, thus excluding them from the fusion process. Concurrently, this approach should dynamically adjust the weights of each client’s model, prioritizing those of higher quality to enhance the global model’s overall accuracy.
Adopting a single-agent reinforcement learning strategy to tackle these issues introduces new challenges. As the number of clients increases, so too does the agent’s action space, prolonging the training duration. Additionally, a single-agent framework is limited to interacting with just one environment, further extending the sampling period. We propose a bifurcated solution inspired by hierarchical reinforcement learning to mitigate these concerns, thereby streamlining the lengthy reinforcement learning training process. This solution comprises two primary stages: the selection of trustworthy clients and the assignment of optimal weights.
Figure 1: The Process of FedDRL framework
Stage 1: During this phase, the objective is to identify K trustworthy models from a pool of N for inclusion in the global model fusion. Identifying clients who have uploaded malicious models is challenging. We address this by employing reinforcement learning to dynamically select and autonomously screen client models, as delineated in equation 2.
$\{M_{a},M_{\mathrm{b}},\ldots,M_{\mathrm{k}}\}\leftarrow\text{SelectTrustworthyModel}\left(\{M_{1},M_{2},\ldots,M_{\mathrm{n}}\}\right)$ (2)
@@
Stage 2: Building on the first step, we then allocate optimal weights to the verified models to bolster the global model’s accuracy, formalized in equation 3.
$\{W_{1},W_{2},\ldots,W_{\mathrm{n}}\}\leftarrow\operatorname{AdaptCalculateWeight}\left(\{M_{a},M_{b},\ldots,M_{\mathrm{k}}\}\right)$ (3)
@@
Here, $AdaptCalculateWeight(.)$ signifies a method for adaptive weight computation, and $W_{i}$ represents the optimal computational weight assigned to each client’s output.
3.2 A Trustworthy Federated Learning Approach Based on Staged Reinforcement Learning
To address these challenges, we introduce a trusted federated learning framework anchored in staged reinforcement learning (FedDRL). This framework unfolds across two distinct phases. In the first phase, we propose an adaptive client selection strategy aimed at identifying and selecting trustworthy clients for participation in model fusion. Subsequently, in the second phase, we formulate a model weight assignment algorithm designed to dynamically allocate fusion weight values to models based on the prevailing fusion environment. The process is depicted in Figure 1.
3.2.1 Adaptive client selection method